OpenStudy (anonymous):

10+x-(2-3x)=4 10+3+x-(2-3+3)=4+3 13+x-(2-6)=7 Is that right how to solve this equation?

OpenStudy (anonymous):

You need to get rid of the parentheses as the first step. Use the distributive property to distribute the minus sign through the parentheses.

OpenStudy (anonymous):

So 10+3+x-2+3-3+3=4?

OpenStudy (anonymous):

\[\begin{align} 10+x-(2-3x)=4\\ 10+x+-1(2-3x)=4\\ 10+x+(-1\cdot2)+(-1\cdot-3x)=4\\ 10+x-2+3x=4 \end{align}\]